# Quantum Kernels for Symbolic Sequences Classification

### v1

@@ 3.2 Structural Generative Kernels
The key idea is these kernels to estimate the similarity between time series by matching of their substructures. The structure of each time series is defined by the evolution of its hidden quantum state. Therefore it is reasonable to estimate the structural similarity by the divergence of appropriate statistics of their generative processes. The other intuition supporting the idea to compare the generative statistics is that the observable sequences depend on causal latent factors, represented by the hidden states. We can assume that the semantics of each sequence is defined by the level of participation of each of the hidden states. Therefore the structural similarity can be measured by the divergence of the expectations of the hidden states.
To formalize these ideas, let’s consider the process of generation of a sequence $\mathbf{y}\in L^{Q},\mathbf{y}=y_{1}y_{2}\cdots y_{n},y_{i}\in\Sigma,i\in[1,n]$. Initially the model is in state $\rho_{0}$. After the generation of the first symbol $y_{1}$ the system is state $\rho_{1}$ defined as follows:
$\rho_{1}=\frac{T_{y_{1}}\rho_{0}}{tr(T_{y_{1}}\rho_{0})}$
The state after the generation of the second symbol $y_{2}$ is defined as:
$\rho_{2}=\frac{T_{y_{2}}\rho_{1}}{tr(T_{y_{2}}\rho_{1})}$
Finally, the end state is
$\rho_{n}=\frac{T_{y_{n}}\rho_{n-1}}{tr(T_{y_{n}}\rho_{n-1})}$
We propose to map the sequence of symbols $\mathbf{y}=y_{1}y_{2}\cdots y_{n}$ to the expected state of the its generative sequence of states $\rho_{1}\rho_{2}\cdots\rho_{n}$ as follows:
$\phi_{s}(\mathbf{y})=\hat{\rho}_{\mathbf{y}}=\frac{1}{n}\sum_{i\in[1,n]}\rho_{i}$
It is easy to demonstrate, that the average of a set of density matrices is a density matrix. Therefore, the expectation $\hat{\rho}$ is a quantum state and we can define a valid structural quantum kernel using the trace distance as follows:
@@
A valid structural kernel can be defined using Bures metric as well:
@@ 4 Quantum Kernels in Quantum Circuits Computing Model
In many cases the class of a sequence depends on the current state and the probability of its continuations. For example, in a financial time series of price movements the class of a sequence can be 1 if the next movement is expected to be ”UP”. In an English sentence the class of a phrase is ”subject” if the expected next word is a verb. In these cases we classify the sequence by the distributions of their continuations. Therefore it is natural to assume that if the induced future distributions of two sequences are close, the similarity of these sequences is high. Let’s formalize this intuitive explanation. Let $\mathbf{a}=a_{1}\ldots a_{k}$ and $\mathbf{b}=b_{1}\ldots b_{l}$ are the sequences we want to compare. The model $Q$ defines the following end states for each of them:
@@
Figure 3: Implementation of projective kernel for market QHMM circuit
@@
5 Examples: The Limit Order Book
5.1 Artificial market example with four hidden states
In our previous paper [53] we have discussed the artificial $4$ states market example with binary observable, where the distribution of sequences was defined through classical hidden Markov model with defined state transitions and symbol emissions probability matrices. We showed that $\mathcal{H}_{S}$ (state system) and $\mathcal{H}_{E}$ (emission) are Hilbert spaces with dimension $2$, so the combined space $\mathcal{H}_{E}\otimes\mathcal{H}_{S}$ circuit required only two qubits. QHMM defining market distribution was trained with evolutionary algorithm.
In this section we will construct a kernel matrix for all pairs of sequences. One possible approach to implement this on the hardware is to use SWAP test [54] as shown on FIG 1. Qubits $q_{0}$, $q_{1}$ and $q_{3}$, $q_{4}$ on FIG 1 implement two QHMMs, $q_{2}$ is used for fidelity calculation. Maximally mixed states are prepared for both QHMMs leveraging classical registers $1$ and $3$. Then the circuit implements two steps of QHMM generator. Classical registers $1$ and 00 capture the symbols of the first sequence, registers $3$ and $2$ of the second. Finally, the probability of measuring 00 on qubit $q_{2}$ captured on classical register $4$ can be converted to the squared fidelity [54] as
$|\langle{\phi}|\psi\rangle|^{2}=2\times(-0.5+\#\{(0)\}_{q_{2}}\times R^{-1})$
where $\phi$ and $\psi$ are statevectors of respective state systems of both QHMMs, #$\#\{(0)\}_{q_{2}}$ is the number of observed $(0)$ on qubit $q_{2}$ and $R$ is the number of shots. The number of shots required is usually estimated as $O(n)$, where $n$ is the number of states. In our case, $n=2^{2t+1}$, where $t$ is the length of the sequence. Unfortunately, for sequences of length 7-8 this requires about 1 million shots or more. Another downside is that this approach assumes calculating kernel element for pure states. In general, we can be working with mixed states. A better approach is to use projective kernels [55]. It is well-known [39, 56] that a density matrix of a single qubit can be written as
@@ 5.1 Artificial market example with four hidden states
[…]
@@ 6 Conclusions and Outlook
Discussed quantum computing approach to the design of similarity measures and kernels for classification of stochastic symbol time series. Similarity of the states implies (probabilistic) similarity of the distributions defined by the states and the processes originating from these states. This is the heuristic we use in order to propose this class of kernels for classification of sequences, based on their future behavior. The proposed approach can be applied for classification of high frequency symbolic time series in the financial industry. We have shown two different approaches to calculating temporal quantum kernels on quantum computers and suggested multi-programming approach to increase utilization of quantum resources. We have shown that our approach leads to good separation between classes.
7 Acknowledgements

### v2

@@ 4.2 Structural Generative Kernels
The classification maps of the structural classification tasks (11) depend on the probabilistic process that generates the sequences. The generative process is defined by the evolution of model’s quantum state. For any example $\mathbf{y}\in L^{Q},\mathbf{y}=y_{1}y_{2}\cdots y_{n},y_{i}\in\Sigma,i\in[1,n]$ a generating state sequence of $\mathbf{y}$ is defined as (23):
$\left\{\rho_{i}=\frac{T_{y_{i}}\rho_{i-1}}{tr(T_{y_{i}}\rho_{i-1})}:i\in[1,n]\right\}.$ (46)
We introduce a kernel that estimates the structural similarity of examples based on the divergence of appropriate statistics derived from the generating state sequences. The structure of each sequence depends on the level of involvement of model’s state in the generative process: if a state participates more often, its impact is stronger. To account for this dependency, the structural kernel maps each example to the expectation of its generating states. Then, the structural similarity of sequences is measured by the divergence of the expectations of their generating states. The map of examples to the expectations of quantum states is defined using (46) as follows:
$\phi_{s}(\mathbf{y})=\hat{\rho}_{\mathbf{y}}=\frac{1}{n}\sum_{i\in[1,n]}\rho_{i},$
where $\mathbf{y}=y_{1}y_{2}\cdots y_{n}$. It is easy to demonstrate, that the average of a set of density matrices is a density matrix. Therefore, the expectation $\hat{\rho}$ is a quantum state and we can define a valid structural quantum kernel using the trace distance as follows:
@@
A valid structural kernel can also be defined using the Bures metric:
@@
5 Empirical Evaluation of Quantum Generative Kernels
In this section, we use the stochastic process language discussed in Example 1 and the corresponding generative QHMM (Figure 3) to empirically investigate the behaviour of the introduced kernels in several contexts: dimension of the quantum Hilbert space, type of the classification task, and type of the stochastic divergence measures of the quantum states. The QHMM discussed in Example 1 is minimal, with dimension of the state Hilbert space $N=2$ (2). To study the impact of the size of the quantum Hilbert space on kernels behaviours we learned and investigated two larger models of the same stochastic process language with dimensions $N=4$ and $N=16$ correspondingly. The components of these models: initial state preparation, transition algorithm, and observable emission in circuit quantum computing model are presented in appendix A.
5.1 Quantum Feature Space Dimension Impact
It is important to demonstrate how the introduced kernels support the expectation that mapping the original domain into a high-dimensional Hilbert space enhances the separability of examples. For QHMMs of the Market example with Hilbert space dimensions 2, 4, and 16, we measured the distributions of the distances between examples induced by the kernels. The results are presented in Figures 5 and 6. If a kernel maps a significant number of pairs to short distance ranges, indicating they are not well distinguishable, this behavior could pose a potential problem for kernel learning methods to perform effectively in nonlinear environments. On the other hand, if a kernel assigns significant distances to most pairs of examples, it might be more successful in challenging learning tasks.
Figure 5: Number of Pairs by Distance - Trace Metric
Figure 6: Number of Pairs by Distance - Bures Metric
The initial observation reveals that the Predictive kernels in low-dimensional Hilbert spaces struggle to distinguish half of the examples while effectively separating the other half. On the other hand, the Structural kernels on 1 and 2 Qubits display better performance by effectively separating a significant portion of the examples. Furthermore, it is evident that the Trace metric yields better separation than the Bures metric. Notably, the 4 Qubit kernel consistently performs well across all scenarios, as the Structural kernel with the Trace metric achieves perfect separation of all pairs of examples. This behavior underscores the direct impact of increasing the dimension of the Hilbert space of the QHMM on the kernel’s performance.
5.2 Kernel-Induced Distances with Respect to Classification Tasks
We demonstrate that distances calculated by the structural and predictive kernels are correlated with the classes of corresponding examples. Specifically, examples at shorter distances tend to belong to the same class. We will consider simple structural and predictive classification tasks as defined in appendix B. On figures 7 and 8 it is evident that the distances defined by the 4-qubit kernels exhibit an inverse linear relationship with the probability of examples belonging to the same class. The stronger correlation is demonstrated by the 4-qubit predictive kernel in the predictive classification task. Conversely, the performance of the 1-qubit and 2-qubit structural and predictive kernels is the poorest in the structural classification task. Similar results, presented in appendix C, were obtained for the trace metric. Ultimately, the dimensionality of the quantum feature space significantly impacts kernels’ performance.
Figure 7: Probability of Examples Belonging to the Same Class - Predictive Kernel, Bures Metric
Figure 8: Probability of Examples Belonging to the Same Class - Structural Kernel, Bures Metric
5.3 Quantum Kernels vs Classical Kernels
To compare the performance of the proposed kernels against classical ones, we use the classification tasks described in appendix B. Three common algorithms [52] - random forest classifier, support vector classifier, and kernelized k-nearest neighbours classifier were applied to the classification tasks. The random forest classifier was used as a bench mark and the other were implemented with classical and quantum kernels. As classical kernel was the radial basis function (RBF) kernel defined as follows [14]:
$K(\mathbf{y}^{1},\mathbf{y}^{2})=\operatorname{exp}\left(-\frac{\left\lVert(\mathbf{y}^{1}-\mathbf{y}^{2})\right\rVert^{2}}{2\sigma^{2}}\right),$
where $\left\lVert(\mathbf{y}^{1}-\mathbf{y}^{2})\right\rVert$ is the Euclidean distance between the sequences, and $\sigma$ is a parameter. In the quantum implementation of the RBF kernel we replaced the the Euclidean distance with the quantum predictive kernel $\kappa_{p}(\mathbf{y}^{1},\mathbf{y}^{2})$ (37),(45):
$K_{q}(\mathbf{y}^{1},\mathbf{y}^{2})=\operatorname{exp}\left(-\frac{\kappa_{p}(\mathbf{y}^{1},\mathbf{y}^{2})}{2\sigma^{2}}\right),$
and the quantum structural kernel $\kappa_{s}(\mathbf{y}^{1},\mathbf{y}^{2})$ (47), (48):
$K_{q}(\mathbf{y}^{1},\mathbf{y}^{2})=\operatorname{exp}\left(-\frac{\kappa_{s}(\mathbf{y}^{1},\mathbf{y}^{2})}{2\sigma^{2}}\right),$
In all structural and predictive task scenarios, the quantum kernels exhibited superior performance compared to their classical counterparts and the RFS benchmark. The algorithms were tested on 500 samples split to training and testing at $50\%$ using Trace metric (37) for the kernels. As it can be seen in Table 2 and Table 3 in both classification tasks the quantum kernels outperformed their classical counterparts as the SVC with quantum kernel outperformed the RFS benchmark algorithm as well.
Similar performance were observed with kernels using Bures metric (45), (48) as the results are presented in appendix D.
Classifier Kernel In Sample CI Out Sample CI
RFS N/A 0.998 0.996 - 0.999 0.968 0.966 - 0.971
SVC Classical 0.951 0.947 - 0.955 0.916 0.910 - 0.922
SVC Quantum 0.999 0.999 - 1.000 0.980 0.977 - 0.981
k-NN Classical 0.971 0.969 - 0.973 0.916 0.913 - 0.919
k-NN Quantum 0.993 0.992 - 0.994 0.980 0.978 - 0.982
Table 2: Kernels Performance by Recognizer Accuracy: Predictive Task, Predictive Kernel, Distance Metric- Trace, CI=95%, Kernel Type RBF
Classifier Kernel In Sample CI Out Sample CI
RFS N/A 0.985 0.984 - 0.987 0.824 0.819 - 0.829
SVC Classical 0.962 0.959 - 0.965 0.903 0.896 - 0.909
SVC Quantum 1.000 - 0.949 0.946 - 0.952
k-NN Classical 0.863 0.857 - 0.869 0.720 0.715 - 0.723
k-NN Quantum 0.950 0.948 - 0.953 0.893 0.889 - 0.897
Table 3: Kernels Performance by Recognizer Accuracy: Structural Task, Structural Kernel, Distance Metric- Trace, CI=95%, Kernel Type RBF
@@ 6 Quantum Kernels in Quantum Circuits Computing Model
In many cases the class of a sequence depends on the current state and the probability of its continuations. For example, in a financial time series of price movements the class of a sequence can be 1 if the next movement is expected to be “UP”. In an English sentence the class of a phrase is “subject” if the expected next word is a verb. In these cases we classify the sequence by the distributions of their continuations. Therefore it is natural to assume that if the induced future distributions of two sequences are close, the similarity of these sequences is high. Let’s formalize this intuitive explanation. Let $\mathbf{a}=a_{1}\ldots a_{k}$ and $\mathbf{b}=b_{1}\ldots b_{l}$ are the sequences we want to compare. The model $Q$ defines the following end states for each of them:
@@
Figure 11: Implementation of projected kernel for market QHMM circuit
@@
In this section we will construct a kernel matrix for all pairs of sequences. One possible approach to implement this on the hardware is to use SWAP test [65] as shown on FIG 9. Qubits $q_{0}$, $q_{1}$ and $q_{3}$, $q_{4}$ on FIG 9 implement two QHMMs, $q_{2}$ is used for fidelity calculation. Maximally mixed states are prepared for both QHMMs leveraging classical registers $1$ and $3$. Then the circuit implements two steps of QHMM generator. Classical registers $1$ and 00 capture the symbols of the first sequence, registers $3$ and $2$ of the second. Finally, the probability of measuring 00 on qubit $q_{2}$ captured on classical register $4$ can be converted to the squared fidelity [65, 66] as
$|\langle{\phi}|\psi\rangle|^{2}=2\times(-0.5+r_{0}\times R^{-1})$
where $\phi$ and $\psi$ are state vectors of respective state systems of both QHMMs, $r_{0}$ is the number of observed zero bit-strings on qubit $q_{2}$ and $R$ is the number of shots. The number of shots required is usually estimated as $O(n)$, where $n$ is the number of states. In our case, $n=2^{2t+1}$, where $t$ is the length of the sequence. Unfortunately, for sequences of length 7-8 this requires about 1 million shots or more. Another downside is that this approach assumes calculating kernel element for pure states. In general, we can be working with mixed states. A better approach is to use projected kernels [67]. It is well-known [50, 68] that a density matrix of a single qubit can be written as
@@
The circuit implementation [68] is shown on FIG 10. We can implement projected kernel for the market QHMM circuit as shown on FIG 11. Finally, the kernel element is constructed from two density matrices as $k_{1,2}=e^{-\gamma\lVert\rho_{1}-\rho_{2}\rVert^{2}}$, where $\lVert$.$\rVert$ . ∥ is a Frobenius norm and $\gamma$ is a hyperparameter (for now, we set $\gamma=1$). In order to boost the effective number of shots we propose to run multiple circuits simultaneously in parallel on a single chip by combining them on a single circuit (multi-programming [69, 70, 71]). With this in mind, the initial qubit layout on IBM Nazca device is shown on FIG 13, where we effectively use $72$ qubits. Color bar plot of the kernel matrix is shown on FIG 13. We see a perfect separation between classes. In this case the main system required only $1$ qubit. In case, we have systems of many qubits we can use 1-reduced density matrix (1-RDM) or N-RDM approximations.
@@ 7 Conclusions and Outlook
In this article, we introduced a generative quantum computing approach to designing similarity measures and associated kernels for the classification of stochastic symbolic time series. The proposed kernels are built upon a novel class of quantum generative models known as quantum hidden Markov models. We provided theoretical justifications of the efficacy of the kernels in predictive classification tasks. Extensive simulation experiments have demonstrated the scalable and discriminative performance of the kernels in high-dimensional Hilbert spaces. Furthermore, we have devised and implemented the kernels in the quantum circuits computing model and successfully conducted experiments on quantum hardware. Finally, we compared the performance of the introduced quantum kernels to their classical counterparts on simple predictive and structural classification tasks, where the quantum kernels showed clear superiority. The proposed approach introduces novel quantum generative modeling techniques for designing hybrid classical-quantum algorithms for clustering, classification, and regression tasks in the domain of symbol sequences and time series. An active research direction involves applying the proposed kernels to the classification of high-frequency time series in the field of quantitative finance. Success in this challenging domain will require extending the underlying quantum stochastic model to encompass the generation of fractional and other non-Markovian stochastic processes.
Acknowledgements
@@
Appendix A Market Quantum Generative Models in Higher Dimensions
Figure 15: Market Movements QHMM: Two State Qubits; One Emission Qubit
Figure 16: Market Movements QHMM: 4 State Qubits; One Emission Qubit
Appendix B Classification Tasks
We empirically study binary classification problem for binary time series in the domain for market movement process (Example 1). The generative model of the example domain is a QHMM (26) parameterized in Hilbert spaces with dimensions 2, 4 and 8 (Figures 3, 15, 16).
A structural classificataion task is defined by the following class-mapping function (11):
[…]
The predictive classification task uses class-mapping function (12) defined for forward sequences with length $k=5$:
$p_{p}(\mathbf{y},k,c)=L[c_{2}],$
where $c_{2}$ is the binary number corresponding to the prefix and the label of each prefix is defined by
$L=^{\prime}11110011001100100110001011000110^{\prime}$
The examples $\mathbf{y}\in\{0,1\}^{*}$ are sampled with probability (22):
$P\bigl{[}\mathbf{y}|\rho_{0}\bigr{]}=\operatorname{tr}(T_{\mathbf{y}}\rho_{0}).$
Appendix C Kernel Induced Distances Across Classification Tasks - Trace Metric
Figure 17: Probability of Examples Belonging to the Same Class -Predictive Kernel, Trace Metric
Figure 18: Probability of Examples Belonging to the Same Class - Structural Kernel, Trace Metric
Appendix D Kernels Performance - Distance Metric Bures
Classifier Kernel In Sample CI Out Sample CI
RFS N/A 0.985 0.984 - 0.987 0.824 0.819 - 0.829
SVC Classical 0.962 0.959 - 0.965 0.903 0.896 - 0.909
SVC Quantum 0.998 0.997 - 0.998 0.947 0.944 - 0.949
k-NN Classical 0.863 0.857 - 0.869 0.720 0.715 - 0.723
k-NN Quantum 0.953 0.951 - 0.956 0.894 0.891 - 0.898
Table 4: Kernels Performance by Recognizer Accuracy: Structural Task, Structural Kernel, Distance Metric- Bures, CI=95%, Kernel Type RBF
Classifier Kernel In Sample CI Out Sample CI
RFS N/A 0.998 0.996 - 0.999 0.968 0.966 - 0.971
SVC Classical 0.951 0.947 - 0.955 0.916 0.910 - 0.922
SVC Quantum 0.999 0.998 - 0.999 0.977 0.975 - 0.979
k-NN Classical 0.971 0.969 - 0.973 0.916 0.913 - 0.919
k-NN Quantum 0.994 0.993 - 0.995 0.979 0.977 - 0.981
Table 5: Kernels Performance by Recognizer Accuracy: Predictive Task, Predictive Kernel, Distance Metric- Bures, CI=95%, Kernel Type RBF